OpenStudy (anonymous):

Choose the equation of the line that contains the points (-3, 2) and (5, -5).

OpenStudy (amistre64):

(-3,2) -(5,-5) ------ -8,7 ; slope = y/x = -7/8

OpenStudy (amistre64):

y = (-7/8)(x+3)+2 y = (-7/8)x -21/8 + 16/8 y = (-7/8)x - 5/8